# Einstein from Noise: Statistical Analysis

### v2

@@ Appendix D High-dimensional signal convergence
Proposition D.1.
Let $\mathsf{S}\sim{\cal N}(\mathsf{\mu},\mathsf{\Sigma})$ be a $d$-dimensional Gaussian random vector, with mean $\mathsf{\mu}$ and a covariance matrix $\mathsf{\Sigma}$. Assume that $\left|\mathsf{\Sigma}_{ij}\right|=\rho_{\left|{i-j}\right|}$, where $\{\rho_{\ell}\}_{\ell\in\mathbb{N}}$ is a sequence of real-valued numbers such that $\rho_{0}=1$, $\rho_{d}<1$, and $\rho_{\ell}\log\ell\rightarrow 0$, as $\ell\to\infty$. Assume also that $\sqrt{\log d}\cdot\max_{1\leq i\leq d}{\left|\mu_{i}\right|}\rightarrow 0$, as $d\rightarrow\infty$, and let $\hat{\mathsf{R}}\triangleq\operatorname*{arg\,max}{\left\{\mathsf{S}_{0},\mathsf{S}_{1},\ldots,\mathsf{S}_{d-1}\right\}}$. Then, for a bounded deterministic function $f:\{0,1,\ldots,d-1\}\rightarrow\mathbb{R}$, we have,
$\lim_{d\to\infty}\mathbb{E}[{f(\hat{\mathsf{R}})}]-\frac{\sum_{r=0}^{d-1}f(r)e^{\mu_{r}a_{d}}}{\sum_{r=0}^{d-1}e^{\mu_{r}a_{d}}}=0,$ (D.1)
@@
The proof of Proposition D.1 is based on an auxiliary result, which we prove in Section D.1. To state this result, we introduce some additional notation. Let $\mathsf{S}(r)$, for $r\in\{0,1,\ldots,d-1\}$, be a discrete stochastic process. We define the function $h^{(\alpha)}(r)$ as follows,
$h^{(\alpha)}(r)\triangleq\mathsf{S}(r)+\alpha f(r),$ (D.2)
@@
$\mathsf{M}_{d}(\alpha)\triangleq\max_{r}h^{(\alpha)}(r),$ (D.3)
@@
$\hat{\mathsf{R}}(\alpha)\triangleq\operatorname*{arg\,max}_{r}h^{(\alpha)}(r).$ (D.4)
Note that $\mathsf{M}_{d}(a)$ and $\hat{\mathsf{R}}(a)$ are random variables. Finally, we denote $\hat{\mathsf{R}}\triangleq\hat{\mathsf{R}}(0)$. We have the following result, which is proved in Appendix D.1.
Lemma D.2.
@@
$\mathbb{E}[f(\hat{\mathsf{R}})]=\left.\frac{\mathrm{d}}{\mathrm{d}\alpha}\mathbb{E}[\mathsf{M}_{d}(\alpha)]\right|_{\alpha=0}.$ (D.5)
Lemma D.2 implies that finding the expected value of $f(\hat{\mathsf{R}})$ is related directly to the derivative of the expected value of the maximum around zero. Thus, the problem of finding the expected value of $f(\hat{\mathsf{R}})$ is related to finding the expected value of the maximum of the stochastic process. In our case, $\mathsf{S}$ will be a Gaussian vector with mean given by (A.19) and a covariance matrix given by (A.20). Thus, our goal now is to find the expected value of the maximum of $\mathsf{S}$. For this purpose, we will recall some well-known results on the maximum of Gaussian processes.
@@
$\mathbb{P}\left[a_{n}(\mathsf{M}_{n}-b_{n})\leq x\right]\to e^{-e^{(-x)}},$ (D.6)
@@
$b_{n}\triangleq\sqrt{2\log n}-\frac{1}{2}\frac{\log{\log{n}}+\log{4\pi}}{\sqrt{2\log n}}.$ (D.8)
@@
It turns out that the above convergence result remains valid even if the sequence $\left\{{\xi_{n}}\right\}$ is not independent and normally distributed. Specifically, as shown in [26, Theorem 6.2.1], a similar result holds for Gaussian random variables $\left\{{\xi_{n}}\right\}$ with a covariance matrix that decays such that $\lim_{n\to\infty}\rho_{n}\cdot\log{n}=0$, and with a mean vector whose maximum value decays faster than $\lim_{n\to\infty}\max_{0\leq m\leq n-1}\left|\mu_{m}\right|\cdot\sqrt{\log{n}}=0$. These conditions precisely match those specified in Theorem 4.3.
Proof of Proposition D.1.
$\mathsf{S}|\mathsf{N}[k]=\mathsf{Z}+\mathsf{\mu},$ (D.9)
$h^{(\alpha)}(r)\triangleq\mathsf{Z}(r)+\mathsf{\mu}(r)+\alpha f(r),$ (D.10)
where we use the same notations as in Lemma D.2. Then, using Lemma D.2,
$\mathbb{E}[f(\hat{\mathsf{R}})]=\left.\frac{\mathrm{d}}{\mathrm{d}\alpha}\mathbb{E}\mathsf{M}_{d}(\alpha)\right|_{\alpha=0},$ (D.11)
Using [26, Theorem 6.2.1], under the assumptions of Proposition D.1, for a sufficiently small value of $\alpha$ such that $\lim_{d\to\infty}\left|\alpha\right|\max_{r}{\left|f(r)\right|}\cdot\sqrt{\log d}=0$, we have for any $x\geq 0$,
$\lim_{d\to\infty}\mathbb{P}\left[a_{d}(\mathsf{M}_{d}(\alpha)-b_{d}-m_{d}^{\star}(\alpha))\leq x\right]=e^{-e^{(-x)}},$ (D.12)
where $a_{d}$ and $b_{d}$ are given in (D.7) and (D.8), respectively, and
$m_{d}^{\star}(\alpha)\triangleq a_{d}^{-1}\log{\left(d^{-1}\sum_{i=0}^{d-1}{e^{a_{d}(\mu_{i}+\alpha f(i))}}\right)}.$ (D.13)
For brevity, we denote,
$\mathsf{T}_{d}(\alpha)\triangleq a_{d}\cdot[\mathsf{M}_{d}(\alpha)-b_{d}-m_{d}^{\star}(\alpha)],$ (D.14)
and we note that,
$\mathsf{T}_{d}(\alpha)-\mathsf{T}_{d}(0)=a_{d}[(\mathsf{M}_{d}(\alpha)-m_{d}^{\star}(\alpha))-(\mathsf{M}_{d}(0)-m_{d}^{\star}(0))],$ (D.15)
and so,
$\Delta_{d}(\alpha)$ $\triangleq\frac{1}{a_{d}}\frac{\mathsf{T}_{d}(\alpha)-\mathsf{T}_{d}(0)}{\alpha}=\frac{\mathsf{M}_{d}(\alpha)-\mathsf{M}_{d}(0)}{\alpha}-\frac{m_{d}^{\star}(\alpha)-m_{d}^{\star}(0)}{\alpha},$ (D.16)
Lemma D.3.
For any $\alpha\neq 0$,
$\lim_{d\to\infty}\left|\Delta_{d}(\alpha)\right|=0,$ (D.17)
Proof of Lemma D.3.
To prove (D.17), we will first show that $\Delta_{d}(\alpha)$ converges to zero in probability. Because $\Delta_{d}(\alpha)$ is uniformly integrable, this is sufficient for the desired $\mathcal{L}^{1}$ convergence above. Specifically, recall from (D.12) that $\mathsf{T}_{d}(\alpha)$ converges in distribution to the Gumbel random variable $\mathsf{Gum}$ with location zero and unit scale, i.e., $\mathsf{T}_{d}(\alpha)\xrightarrow[]{{\cal D}}\mathsf{Gum}$, as $d\to\infty$. Furthermore, it is clear that $\frac{1}{a_{d}}=\frac{1}{\sqrt{2\log d}}\to 0$, as $d\to\infty$. Thus, Slutsky’s theorem [45] implies that,
$\frac{\mathsf{T}_{d}(\alpha)}{a_{d}}\xrightarrow[]{{\cal D}}0.$ (D.18)
$\frac{\mathsf{T}_{d}(\alpha)}{a_{d}}\xrightarrow[]{{\cal P}}0.$ (D.19)
$\Delta_{d}(\alpha)\xrightarrow[]{{\cal P}}0,$ (D.20)
for every $\alpha\neq 0$.
Next, we show that $\Delta_{d}(\alpha)$ is bounded with probability one. Indeed, by the definition of $\mathsf{M}_{d}(\alpha)$ in (D.3), we have,
$\left|\frac{\mathsf{M}_{d}(\alpha)-\mathsf{M}_{d}(0)}{\alpha}\right|\leq\max_{0\leq r\leq d-1}\left|f(r)\right|<\mathsf{C}<\infty,$ (D.21)
$\frac{\mathrm{d}}{\mathrm{d}\alpha}m_{d}^{\star}(\alpha)=\frac{\sum_{i=0}^{d-1}{f(i)\exp\{a_{d}(\mu_{i}+\alpha f(i))}}{\sum_{i=0}^{d-1}{\exp\{a_{d}(\mu_{i}+\alpha f(i))}\}},$ (D.22)
which is bounded because,
$\left|\frac{\sum_{i=0}^{d-1}{f(i)\exp\{a_{d}(\mu_{i}+\alpha f(i))}}{\sum_{i=0}^{d-1}{\exp\{a_{d}(\mu_{i}+\alpha f(i))}\}}\right|\leq\max_{0\leq r\leq d-1}\left|f(r)\right|<\mathsf{C}<\infty.$ (D.23)
Combining (D.16), (D.21) and (D.23), leads to,
$\left|\Delta_{d}(\alpha)\right|$ $\leq\left|\frac{\mathsf{M}_{d}(\alpha)-\mathsf{M}_{d}(0)}{\alpha}\right|+\left|\frac{m_{d}^{\star}(\alpha)-m_{d}^{\star}(0)}{\alpha}\right|\leq 2\max_{0\leq r\leq d-1}\left|f(r)\right|<\infty.$ (D.24)
Now, since $\Delta_{d}(\alpha)$ is bounded, it is also uniformly integrable, and thus when combined with (D.20) we may conclude that,
$\Delta_{d}(\alpha)\xrightarrow[]{\mathcal{L}^{1}}0,$ (D.25)
as claimed. ∎
We continue with the proof of Proposition D.1. First, we show that,
$\lim_{d\to\infty}\lim_{\alpha\to 0}\mathbb{E}\left[\Delta_{d}(\alpha)\right]=\lim_{\alpha\to 0}\lim_{d\to\infty}\mathbb{E}\left[\Delta_{d}(\alpha)\right].$ (D.26)
Indeed, note that,
$\lim_{d\to\infty}\lim_{\alpha\to 0}\mathbb{E}\left[\Delta_{d}(\alpha)\right]=\lim_{d\to\infty}\lim_{\alpha\to 0}\int{\left[\frac{\mathsf{T}_{d}(\alpha)-\mathsf{T}_{d}(0)}{\alpha}\right]\mathrm{d}\mu},$ (D.27)
where $\mathrm{d}\mu$ is the probability measure associated with $\mathsf{T}_{d}$. From (D.24) we know that $\Delta_{d}(\alpha)$ is bounded. Thus, applying the dominated convergence theorem, we obtain,
$\lim_{d\to\infty}\lim_{\alpha\to 0}\int{\left[\frac{\mathsf{T}_{d}(\alpha)-\mathsf{T}_{d}(0)}{\alpha}\right]\mathrm{d}\mu}=\lim_{d\to\infty}\int{\lim_{\alpha\to 0}\left[\frac{\mathsf{T}_{d}(\alpha)-\mathsf{T}_{d}(0)}{\alpha}\right]\mathrm{d}\mu}.$ (D.28)
Since the integral at the right-hand-side of (D.28) is finite and bounded for each value of $\alpha$, and for each value of $d$, the order of the limits can be exchanged, thus leading to (D.26). Therefore, from (D.16) and (D.26), we have,
$\lim_{\alpha\to 0}\lim_{d\to\infty}\mathbb{E}\left[\Delta_{d}(\alpha)\right]$ $=\lim_{d\to\infty}\lim_{\alpha\to 0}\left[\frac{\mathbb{E}[\mathsf{M}_{d}(\alpha)-\mathsf{M}_{d}(0)]}{\alpha}-\frac{[m_{d}^{\star}(\alpha)-m_{d}^{\star}(0)]}{\alpha}\right]$ (D.29)
$=\lim_{d\to\infty}\left.\left[\frac{\mathrm{d}}{\mathrm{d}\alpha}\mathbb{E}\mathsf{M}_{d}(\alpha)-\frac{\mathrm{d}}{\mathrm{d}\alpha}m_{d}^{\star}(\alpha)\right]\right|_{\alpha=0}.$ (D.30)
Now, Lemma D.3 implies that the left-hand-side of (D.29) nullifies, and thus,
$\lim_{d\to\infty}\left.\left[\frac{\mathrm{d}}{\mathrm{d}\alpha}\mathbb{E}\mathsf{M}_{d}(\alpha)-\frac{\mathrm{d}}{\mathrm{d}\alpha}m_{d}^{\star}(\alpha)\right]\right|_{\alpha=0}=0.$ (D.31)
Finally, combining (D.22) and (D.31), we obtain (D.1), which concludes the proof.
D.1 Proof of Lemma D.2
@@
The proof technique of Lemma D.2 is similar to the technique used in [36, 28], but with a non-trivial adaption to the discrete case. To prove this lemma, we will first establish a deterministic counterpart of (D.5). Specifically, we define,
$h^{(\alpha)}(r)\triangleq X(r)+\alpha f(r),$ (D.32)
@@
$s(\alpha)\triangleq\max_{r}\{h^{(\alpha)}(r)\},$ (D.33)
@@
$Z^{(\alpha)}_{\max}\triangleq\operatorname*{arg\,max}_{r}\{h^{(\alpha)}(r)\}.$ (D.34)
@@
$\left.\frac{\mathrm{d}}{\mathrm{d}\alpha}s(\alpha)\right|_{\alpha=0}=f(Z^{(0)}_{\max}).$ (D.35)
Proof of Lemma D.4.
@@
$=\lim_{\alpha\to 0}\frac{\max_{r}[X(r)+\alpha f(r)]-\max_{r}X(r)}{\alpha}.$ (D.36)
@@
$\max_{r}[X(r)+\alpha f(r)]=X(Z^{(\alpha)}_{\max})+\alpha f(Z^{(\alpha)}_{\max}),$ (D.37)
@@
${Z^{(\alpha)}_{\max}}={Z^{(0)}_{\max}}.$ (D.39)
Combining (D.36)–(D.39) yields,
$\left.\frac{\mathrm{d}}{\mathrm{d}\alpha}s(\alpha)\right|_{\alpha=0}$ $=\lim_{\alpha\rightarrow 0}\frac{X(Z^{(\alpha)}_{\max})+\alpha f(Z^{(\alpha)}_{\max})-X(Z^{(0)}_{\max})}{\alpha}$ (D.40)
$=f{(Z^{(0)}_{\max})},$ (D.41)
@@
We are now in a position to prove Lemma D.2. Similarly to the deterministic case, we define the random function,
$h^{(\alpha)}(r)=\mathsf{S}(r)+\alpha f(r),$ (D.42)
where $\mathsf{S}:\{0,1,\ldots,d-1\}\to\mathbb{R}$ is a discrete stochastic process, and $f$ is a deterministic function. We assume that $\mathsf{S}$ has a continuous probability distribution without any single point with a measure greater than 0. Using Lemma D.4, for each realization of $\mathsf{S}(r)$, such that $\mathsf{S}(r)$ is injective, we have,
$f(\hat{\mathsf{R}})=\left.\frac{\mathrm{d}}{\mathrm{d}\alpha}\ \mathsf{M}_{d}(\alpha)\,\right|_{\alpha=0}.$ (D.43)
Under the assumption above of $\mathsf{S}(r)$, the measure of the set of events that $\mathsf{S}$ is not injective is zero. Therefore, the fact that $\frac{\mathsf{M}_{d}(\alpha)-\mathsf{M}_{d}(0)}{\alpha}$ is bounded (see, (D.21)) and (D.43), imply that,
$\mathbb{E}[f(\hat{\mathsf{R}})]$ $=\int{f(\hat{\mathsf{R}})\mathrm{d}\mu}$
@@
$=\left.\frac{\mathrm{d}}{\mathrm{d}\alpha}\ \mathbb{E}\mathsf{M}_{d}(\alpha)\,\right|_{\alpha=0},$ (D.44)
@@
Appendix E Preliminaries for Theorem 4.3
Remark on notation. In this and next sections, we omit the dependence on $10\leq i\leq M-$ 1, where this is clear from the context, e.g., $\mathsf{N}_{i}=\mathsf{N}$ and $\mathsf{\hat{R}}_{i}=\mathsf{\hat{R}}$.

### v3

@@ Appendix C High-dimensional argmax asymptotics
Proposition C.1 (High-dimensional argmax asymptotics).
Let $\mathsf{S}\sim{\cal N}(\mathsf{\mu},\mathsf{\Sigma})$ be a $d$-dimensional Gaussian random vector, with mean $\mathsf{\mu}$ and a covariance matrix $\mathsf{\Sigma}$. Assume that $\left|\mathsf{\Sigma}_{ij}\right|=\rho_{\left|{i-j}\right|}$, where $\{\rho_{\ell}\}_{\ell\in\mathbb{N}}$ is a sequence of real-valued numbers such that $\rho_{0}=1$, $\rho_{\ell}<1$, and $\rho_{\ell}\log\ell\rightarrow 0$, as $\ell\to\infty$. Assume also that $\sqrt{\log d}\cdot\max_{1\leq i\leq d}{\left|\mu_{i}\right|}\rightarrow 0$, as $d\rightarrow\infty$, and let $\hat{\mathsf{R}}\triangleq\operatorname*{arg\,max}{\left\{\mathsf{S}_{0},\mathsf{S}_{1},\ldots,\mathsf{S}_{d-1}\right\}}$. Then, for a bounded deterministic function $f:\{0,1,\ldots,d-1\}\rightarrow\mathbb{R}$, we have,
$\lim_{d\to\infty}\mathbb{E}[{f(\hat{\mathsf{R}})}]-\frac{\sum_{r=0}^{d-1}f(r)e^{\mu_{r}a_{d}}}{\sum_{r=0}^{d-1}e^{\mu_{r}a_{d}}}=0,$ (C.1)
@@
The proof of Proposition C.1 is based on an auxiliary result, which we prove in Section C.1. To state this result, we introduce some additional notation. Let $\mathsf{S}(r)$, for $r\in\{0,1,\ldots,d-1\}$, be a discrete stochastic process. We define the function $h^{(\alpha)}(r)$ as follows,
$h^{(\alpha)}(r)\triangleq\mathsf{S}(r)+\alpha f(r),$ (C.2)
@@
$\mathsf{M}_{d}(\alpha)\triangleq\max_{r}h^{(\alpha)}(r),$ (C.3)
@@
$\hat{\mathsf{R}}(\alpha)\triangleq\operatorname*{arg\,max}_{r}h^{(\alpha)}(r).$ (C.4)
Note that $\mathsf{M}_{d}(a)$ and $\hat{\mathsf{R}}(a)$ are random variables. Finally, we denote $\hat{\mathsf{R}}\triangleq\hat{\mathsf{R}}(0)$. We have the following result, which is proved in Appendix C.1.
Lemma C.2.
@@
$\mathbb{E}[f(\hat{\mathsf{R}})]=\left.\frac{\mathrm{d}}{\mathrm{d}\alpha}\mathbb{E}[\mathsf{M}_{d}(\alpha)]\right|_{\alpha=0}.$ (C.5)
[…]
@@
$\mathbb{P}\left[a_{n}(\mathsf{M}_{n}-b_{n})\leq x\right]\to e^{-e^{(-x)}},$ (C.6)
@@
$b_{n}\triangleq\sqrt{2\log n}-\frac{1}{2}\frac{\log{\log{n}}+\log{4\pi}}{\sqrt{2\log n}}.$ (C.8)
@@
It turns out that the above convergence result remains valid even if the sequence $\left\{{\xi_{n}}\right\}$ is not independent and normally distributed. Specifically, as shown in [29, Theorem 6.2.1], a similar result holds for Gaussian random variables $\left\{{\xi_{n}}\right\}$ with a covariance matrix that decays such that $\lim_{n\to\infty}\rho_{n}\cdot\log{n}=0$, and with a mean vector whose maximum value decays faster than $\lim_{n\to\infty}\max_{0\leq m\leq n-1}\left|\mu_{m}\right|\cdot\sqrt{\log{n}}=0$. These conditions precisely match those specified in Theorem 4.3.
C.1 Proof of Lemma C.2
@@
The proof technique of Lemma C.2 is similar to the technique used in [39, 31], but with a non-trivial adaption to the discrete case. To prove this lemma, we will first establish a deterministic counterpart of (C.5). Specifically, we define,
$h^{(\alpha)}(r)\triangleq X(r)+\alpha f(r),$ (C.9)
@@
$s(\alpha)\triangleq\max_{r}\{h^{(\alpha)}(r)\},$ (C.10)
@@
$Z^{(\alpha)}_{\max}\triangleq\operatorname*{arg\,max}_{r}\{h^{(\alpha)}(r)\}.$ (C.11)
@@
$\left.\frac{\mathrm{d}}{\mathrm{d}\alpha}s(\alpha)\right|_{\alpha=0}=f(Z^{(0)}_{\max})$. (C.12)
Proof of Lemma C.3.
@@
$=\lim_{\alpha\to 0}\frac{\max_{r}[X(r)+\alpha f(r)]-\max_{r}X(r)}{\alpha}.$ (C.13)
@@
$\max_{r}[X(r)+\alpha f(r)]=X(Z^{(\alpha)}_{\max})+\alpha f(Z^{(\alpha)}_{\max})$, (C.14)
@@
${Z^{(\alpha)}_{\max}}={Z^{(0)}_{\max}}$. (C.16)
Combining (C.13)–(C.16) yields,
$\left.\frac{\mathrm{d}}{\mathrm{d}\alpha}s(\alpha)\right|_{\alpha=0}$ $=\lim_{\alpha\rightarrow 0}\frac{X(Z^{(\alpha)}_{\max})+\alpha f(Z^{(\alpha)}_{\max})-X(Z^{(0)}_{\max})}{\alpha}$ (C.17)
$=f{(Z^{(0)}_{\max})}$, (C.18)
@@
We are now in a position to prove Lemma C.2. Similarly to the deterministic case, we define the random function,
$h^{(\alpha)}(r)=\mathsf{S}(r)+\alpha f(r),$ (C.19)
where $\mathsf{S}:\{0,1,\ldots,d-1\}\to\mathbb{R}$ is a discrete stochastic process, and $f$ is a deterministic function. We assume that $\mathsf{S}$ has a continuous probability distribution without any single point with a measure greater than 0. Using Lemma C.3, for each realization of $\mathsf{S}(r)$, such that $\mathsf{S}(r)$ is injective, we have,
$f(\hat{\mathsf{R}})=\left.\frac{\mathrm{d}}{\mathrm{d}\alpha}\ \mathsf{M}_{d}(\alpha)\,\right|_{\alpha=0}.$ (C.20)
Under the assumption above of $\mathsf{S}(r)$, the measure of the set of events that $\mathsf{S}$ is not injective is zero. Therefore, the fact that $\frac{\mathsf{M}_{d}(\alpha)-\mathsf{M}_{d}(0)}{\alpha}$ is bounded (see, (C.34)) and (C.20), imply that,
[…]
@@
$=\left.\frac{\mathrm{d}}{\mathrm{d}\alpha}\ \mathbb{E}\mathsf{M}_{d}(\alpha)\,\right|_{\alpha=0},$ (C.21)
@@
C.2 Proof of Proposition C.1
$\mathsf{S}|\mathsf{N}[k]=\mathsf{Z}+\mathsf{\mu},$ (C.22)
$h^{(\alpha)}(r)\triangleq\mathsf{Z}(r)+\mathsf{\mu}(r)+\alpha f(r),$ (C.23)
where we use the same notations as in Lemma C.2. Then, using Lemma C.2,
$\mathbb{E}[f(\hat{\mathsf{R}})]=\left.\frac{\mathrm{d}}{\mathrm{d}\alpha}\mathbb{E}\mathsf{M}_{d}(\alpha)\right|_{\alpha=0},$ (C.24)
Using [29, Theorem 6.2.1], under the assumptions of Proposition C.1, for a sufficiently small value of $\alpha$ such that $\lim_{d\to\infty}\left|\alpha\right|\max_{r}{\left|f(r)\right|}\cdot\sqrt{\log d}=0$, we have for any $x\geq 0$,
$\lim_{d\to\infty}\mathbb{P}\left[a_{d}(\mathsf{M}_{d}(\alpha)-b_{d}-m_{d}^{\star}(\alpha))\leq x\right]=e^{-e^{(-x)}},$ (C.25)
where $a_{d}$ and $b_{d}$ are given in (C.7) and (C.8), respectively, and
$m_{d}^{\star}(\alpha)\triangleq a_{d}^{-1}\log{\left(d^{-1}\sum_{i=0}^{d-1}{e^{a_{d}(\mu_{i}+\alpha f(i))}}\right)}.$ (C.26)
For brevity, we denote,
$\mathsf{T}_{d}(\alpha)\triangleq a_{d}\cdot[\mathsf{M}_{d}(\alpha)-b_{d}-m_{d}^{\star}(\alpha)],$ (C.27)
and we note that,
$\mathsf{T}_{d}(\alpha)-\mathsf{T}_{d}(0)=a_{d}[(\mathsf{M}_{d}(\alpha)-m_{d}^{\star}(\alpha))-(\mathsf{M}_{d}(0)-m_{d}^{\star}(0))],$ (C.28)
and so,
$\Delta_{d}(\alpha)$ $\triangleq\frac{1}{a_{d}}\frac{\mathsf{T}_{d}(\alpha)-\mathsf{T}_{d}(0)}{\alpha}=\frac{\mathsf{M}_{d}(\alpha)-\mathsf{M}_{d}(0)}{\alpha}-\frac{m_{d}^{\star}(\alpha)-m_{d}^{\star}(0)}{\alpha},$ (C.29)
Lemma C.4.
For any $\alpha\neq 0$,
$\lim_{d\to\infty}\left|\Delta_{d}(\alpha)\right|=0,$ (C.30)
Proof of Lemma C.4.
To prove (C.30), we will first show that $\Delta_{d}(\alpha)$ converges to zero in probability. Because $\Delta_{d}(\alpha)$ is uniformly integrable, this is sufficient for the desired $\mathcal{L}^{1}$ convergence above. Specifically, recall from (C.25) that $\mathsf{T}_{d}(\alpha)$ converges in distribution to the Gumbel random variable $\mathsf{Gum}$ with location zero and unit scale, i.e., $\mathsf{T}_{d}(\alpha)\xrightarrow[]{{\cal D}}\mathsf{Gum}$, as $d\to\infty$. Furthermore, it is clear that $\frac{1}{a_{d}}=\frac{1}{\sqrt{2\log d}}\to 0$, as $d\to\infty$. Thus, Slutsky’s theorem [49] implies that,
$\frac{\mathsf{T}_{d}(\alpha)}{a_{d}}\xrightarrow[]{{\cal D}}0.$ (C.31)
$\frac{\mathsf{T}_{d}(\alpha)}{a_{d}}\xrightarrow[]{{\cal P}}0.$ (C.32)
$\Delta_{d}(\alpha)\xrightarrow[]{{\cal P}}0,$ (C.33)
for every $\alpha\neq 0$.
Next, we show that $\Delta_{d}(\alpha)$ is bounded with probability one. Indeed, by the definition of $\mathsf{M}_{d}(\alpha)$ in (C.3), we have,
$\left|\frac{\mathsf{M}_{d}(\alpha)-\mathsf{M}_{d}(0)}{\alpha}\right|\leq\max_{0\leq r\leq d-1}\left|f(r)\right|<\mathsf{C}<\infty,$ (C.34)
$\frac{\mathrm{d}}{\mathrm{d}\alpha}m_{d}^{\star}(\alpha)=\frac{\sum_{i=0}^{d-1}{f(i)\exp\{a_{d}(\mu_{i}+\alpha f(i))}}{\sum_{i=0}^{d-1}{\exp\{a_{d}(\mu_{i}+\alpha f(i))}\}},$ (C.35)
which is bounded because,
$\left|\frac{\sum_{i=0}^{d-1}{f(i)\exp\{a_{d}(\mu_{i}+\alpha f(i))}}{\sum_{i=0}^{d-1}{\exp\{a_{d}(\mu_{i}+\alpha f(i))}\}}\right|\leq\max_{0\leq r\leq d-1}\left|f(r)\right|<\mathsf{C}<\infty.$ (C.36)
Combining (C.29), (C.34) and (C.36), leads to,
$\left|\Delta_{d}(\alpha)\right|$ $\leq\left|\frac{\mathsf{M}_{d}(\alpha)-\mathsf{M}_{d}(0)}{\alpha}\right|+\left|\frac{m_{d}^{\star}(\alpha)-m_{d}^{\star}(0)}{\alpha}\right|\leq 2\max_{0\leq r\leq d-1}\left|f(r)\right|<\infty.$ (C.37)
Now, since $\Delta_{d}(\alpha)$ is bounded, it is also uniformly integrable, and thus when combined with (C.33) we may conclude that,
$\Delta_{d}(\alpha)\xrightarrow[]{\mathcal{L}^{1}}0,$ (C.38)
as claimed. ∎
We continue with the proof of Proposition C.1. First, we show that,
$\lim_{d\to\infty}\lim_{\alpha\to 0}\mathbb{E}\left[\Delta_{d}(\alpha)\right]=\lim_{\alpha\to 0}\lim_{d\to\infty}\mathbb{E}\left[\Delta_{d}(\alpha)\right].$ (C.39)
Indeed, note that,
$\lim_{d\to\infty}\lim_{\alpha\to 0}\mathbb{E}\left[\Delta_{d}(\alpha)\right]=\lim_{d\to\infty}\lim_{\alpha\to 0}\int{\left[\frac{\mathsf{T}_{d}(\alpha)-\mathsf{T}_{d}(0)}{\alpha}\right]\mathrm{d}\mu},$ (C.40)
where $\mathrm{d}\mu$ is the probability measure associated with $\mathsf{T}_{d}$. From (C.37) we know that $\Delta_{d}(\alpha)$ is bounded. Thus, applying the dominated convergence theorem, we obtain,
$\lim_{d\to\infty}\lim_{\alpha\to 0}\int{\left[\frac{\mathsf{T}_{d}(\alpha)-\mathsf{T}_{d}(0)}{\alpha}\right]\mathrm{d}\mu}=\lim_{d\to\infty}\int{\lim_{\alpha\to 0}\left[\frac{\mathsf{T}_{d}(\alpha)-\mathsf{T}_{d}(0)}{\alpha}\right]\mathrm{d}\mu}.$ (C.41)
Since the integral at the right-hand-side of (C.41) is finite and bounded for each value of $\alpha$, and for each value of $d$, the order of the limits can be exchanged, thus leading to (C.39). Therefore, from (C.29) and (C.39), we have,
$\lim_{\alpha\to 0}\lim_{d\to\infty}\mathbb{E}\left[\Delta_{d}(\alpha)\right]$ $=\lim_{d\to\infty}\lim_{\alpha\to 0}\left[\frac{\mathbb{E}[\mathsf{M}_{d}(\alpha)-\mathsf{M}_{d}(0)]}{\alpha}-\frac{[m_{d}^{\star}(\alpha)-m_{d}^{\star}(0)]}{\alpha}\right]$ (C.42)
$=\lim_{d\to\infty}\left.\left[\frac{\mathrm{d}}{\mathrm{d}\alpha}\mathbb{E}\mathsf{M}_{d}(\alpha)-\frac{\mathrm{d}}{\mathrm{d}\alpha}m_{d}^{\star}(\alpha)\right]\right|_{\alpha=0}.$ (C.43)
Now, Lemma C.4 implies that the left-hand-side of (C.42) nullifies, and thus,
$\lim_{d\to\infty}\left.\left[\frac{\mathrm{d}}{\mathrm{d}\alpha}\mathbb{E}\mathsf{M}_{d}(\alpha)-\frac{\mathrm{d}}{\mathrm{d}\alpha}m_{d}^{\star}(\alpha)\right]\right|_{\alpha=0}=0.$ (C.44)
Finally, combining (C.35) and (C.44), we obtain (C.1), which concludes the proof.
Appendix D Proof of Theorem 4.3
Remark on notation. In this section, we omit the dependence on $0\leq i\leq M-1$, where this is clear from the context, e.g., $\mathsf{N}_{i}=\mathsf{N}$ and $\mathsf{\hat{R}}_{i}=\mathsf{\hat{R}}$. In addition, to streamline notation we often omit the explicit dependence on the signal length and write $x$ (and the associated quantities) in place of $x^{(d)}$ whenever the dimension $d$ is understood from the context.
D.1 Notations and auxiliary results